OpenStudy (anonymous):

Find P(X > 1). P(x) 0.30 0.40 0.15 ?

OpenStudy (mathmate):

Is it p(1)=0.3, p(2)=0.4, p(3)=0.15, and p(4)=0.15?

OpenStudy (anonymous):

No its P(x) 1=0.30 2=0.40 3=0.15 4=? and it says find Find P(X > 1).

OpenStudy (mathmate):

Well, sum(P(X))=1, so p(4)=0.15 as I suggested above. P(X>1) = P(X=2)+P(X=3)+P(X=4) = 0.7

OpenStudy (anonymous):

Thanks! I might have more questions.. will you get back to me please?

OpenStudy (mathmate):

If you insist, you can type @mathmate when you type your question. I am logging off in the next few minutes. However, there are many capable mathematicians here, so no worries!